Use calculus to find the MPK and MPL

Find the Slope of the Isoquant

Production Function: Q=50+6L-.5L^2+22K-.25K^2
Answer #1

Q = 50 + 6L -0.5L2 + 22K - 0.25K2

Marginal product of capital , MPK = dQ/dK = 22K1-1 - (2)(0.25)K2-1

= 22 - 0.5K

Marginal product of labor , MPL = dQ/dL = 6L1-1- (2)(0.5)L2-1

= 6 - L

Slope of the isoquant = MRTS = -MPL/MPK = -[(6-L)/(22-0.5K)]